# Multimode Entangled Squeezed Light Generation and Propagation in a Coupled-Cavity Photonic Crystal

### v2

@@ Abstract
We present an efficient and accurate method for modeling the generation and propagation of quantum states of light in lossy coupled-cavity systems. We apply our approach to the design and modeling of a multimode photonic crystal coupled-cavity system for the generation of entangled squeezed states of light on-chip. The system consists of a three-mode, three-defect resonant structure coupled to three coupled-resonator optical waveguides (CROWs) in a square lattice silicon photonic crystal slab. We model the system using a basis of 184 non-orthogonal lossy quasi-modes. A two-mode squeezed thermal state of light is generated via degenerate spontaneous four-wave mixing in the resonant structure, which is pumped with a Gaussian pulse via a pump-CROW. The generated entangled signal and idler pulses of light are coupled out of the resonant structure into the two output-CROWs. Due to careful design of the resonant structure and the CROWs, the signal and idler light in the two output CROWs remains entangled even after propagating tens of cavities down the CROWs. Our approach is general and computationally efficient and thus can be applied to modeling the generation and propagation of quantum states of light in a wide variety of coupled-cavity photonic structures.
@@ I Introduction
Entangled squeezed thermal states of light [1, 2] are sought for their entanglement and noise properties. They lend themselves to applications such as continuous-variable quantum computing [3, 4, 5], quantum key distribution [6], quantum-enhanced sensing [7, 8, 9], and the generation of non-Gaussian entangled states of light [10]. Consequently, there is considerable interest in developing and modeling nanophotonic platforms for the generation of entangled quantum states of light.
@@
Entangled, multi-mode, squeezed thermal states of light may be created by generating single-mode squeezed thermal states of light in an optical cavity [11, 12] and interfering them using a beam-splitter array to build multi-mode entanglement [13]. Alternatively, they can be generated directly via a nonlinear process described by a Hamiltonian containing terms with an operator of the form $(a^{\dagger}_{1}a^{\dagger}_{2})^{n},\ n\in\mathbb{N}^{>0}$, where $a_{i}^{\dagger}$ is the creation operator for photons in the $i^{th}$ mode.
Nonlinear devices using microring resonators have been explored both experimentally [14, 15, 16, 13] and theoretically [17, 18, 12, 19, 20, 21, 22] for the generation of entangled squeezed states of light. Photonic crystal (PC) coupled cavity systems (CCSs) have also been investigated theoretically [23, 24, 25] for the generation of squeezed states of light.
Microring resonator systems can be fabricated with very high quality factors (Q) [16] and are relatively simple to design, fabricate, and couple to channel waveguides. However, they support a large number of nearly equally spaced modes, which can lead to parasitic nonlinear processes that degrade the properties of the desired quantum state [16, 26]. In contrast, although the quality factors of the resonant modes in PC CCSs are generally an order of magnitude or more smaller than can be attained in microrings, the resonant cavities are much smaller, with cavity sizes on the order of hundreds of nanometers. Moreover, they can be designed such that they only support the modes required for the desired optical processes.
In this work, we examine PC CCSs as an alternative platform to microring resonators (see Fig. 1). We exploit the ability to choose the number and frequency spacing of resonant-structure modes and to tailor the pass-band of coupled resonator optical waveguides (CROWs) to design a PC CCS to generate and propagate two-mode entangled squeezed states of light. In addition, we use this system to demonstrate the effectiveness of our theoretical approach in simulating the generation of multi-photon quantum states of light in complicated dielectric nanostructures.
Any nanophotonic system used to generate entangled squeezed states, regardless of the mechanism of generation, is a complicated dielectric structure with multiple interacting elements. Such composite systems are difficult to model and simulate. Any theoretical framework employed must be be able to model not only each individual element but also their mutual interactions and their loss to the environment. In this work, we present an approach to simulating the nonlinear generation of quantum states of light in systems of coupled lossy nanophotonic resonators in which there is a localized region – the resonant structure (RS) – where the nonlinear interaction occurs. Our approach uses a basis of non-orthogonal, lossy quasi-modes (QMs) [27, 28] to account for the loss characteristic of the open resonators from which the system is constructed. We focus our attention on a PC CCS in this work, but our method can be applied to other nanophotonic resonator systems.
Our theoretical and computational approach involves four distinct steps. First, we determine the QMs of the full system under consideration. To accomplish this, we use finite difference time domain (FDTD) simulations to compute the electric field profile and quality factor of the QMs of the single-defect cavities from which we construct the PC CCS [29]; we then use a tight-binding approach to construct the QMs of the full system from these single-defect modes [29, 27, 28, 30]. Second, we use this basis of system QMs to compute the classical evolution of the pump light (for a given initial pump field) and determine the pumping-field amplitude inside the RS as a function of time. Third, we compute the time evolution of density operator of the quantum state of light generated via nonlinear interaction in the RS. Fourth, we solve the Lindblad master equation after the pump has peaked in order to determine the quantum evolution of the generated state throughout the system.
@@
To demonstrate our approach, we design and simulate a PC CCS capable of generating two-mode entangled squeezed thermal states directly. The system consists of a three-mode, three-defect RS coupled to three CROWs [31] in a square lattice silicon photonic crystal slab (see Fig. 1). The PC CCS we propose is a spatially smaller alternative to the microring resonator-based systems [21, 14, 15, 16, 26]. The entangled two-mode squeezed thermal state is generated in the RS at the signal and idler frequencies via spontaneous four-wave mixing (SFWM) when excited by a Gaussian pump pulse at the pump frequency, which is midway between the signal and idler frequencies. The key element of the system is the RS, which is comprised of three single-defect cavities and supports three resonant modes, one each at the signal, pump, and idler frequencies. The RS enhances the nonlinear spontaneous four wave mixing process that generates the desired entangled state. The generated light propagates out of the RS as two entangled pulses at the signal and idler frequencies in spatially-distinct CROWs that are designed to block any pump light.
@@
The paper is organized as follows: in Section II, we provide a detailed description of the designed device and outline the design goals; in Section III, we introduce the four-step analysis procedure we developed to simulate the designed system; in Section IV, we present and discuss the results of the analysis; and finally in Section V, we present our conclusions.
@@ II System Overview and Design Goals
In this section, we describe the PC CCS that we have designed and optimized using our theoretical approach. We present our design process, outline the objectives of our design, and present the specifications of the designed PC CCS within which we will simulate the generation and propagation of multimode squeezed light.
The structure is a PC CCS with a three-mode RS and three CROWs in a PC slab, shown schematically in Fig. 1. The underlying PC slab is a 384 nm thick silicon (n=3.47) slab with air holes of radius 192 nm etched in a rectangular lattice with a lattice constant $a=480$ nm. This particular square lattice geometry of the PC was chosen because the single-defect cavities formed in this PC support a single relatively high-Q mode in the photonic band-gap of the PC [29].
We have designed the structure to generate and propagate entangled two-mode (2M) squeezed thermal states (STS) with high brightness and robust entanglement. We have four main objectives for the design of the PC CCS. First, we need a RS with spatially-overlapping modes at the frequencies of the pump ($\omega_{P}$), signal ($\omega_{S}$), and idler ($\omega_{I}$), where $2\omega_{p}=\omega_{s}+\omega_{I}$ so that the process of SFWM can result in the generation of a quantum state of light in the signal and idler modes. Second, we need to be able to inject the pump light into the RS via a CROW such that there is a large pump intensity buildup inside of the RS. Third, we require that the quantum state of light generated in the RS can couple out of the RS into two separate CROWs, with central frequencies, $\omega_{s}$ and $\omega_{I}$, that do not allow propagation of the pump light.
@@
The three CROWs were designed such that their bandwidths encompass only one of the pump, signal, or idler frequencies. This is achieved by constructing each CROW from a single-defect cavity with a resonant frequency equal to the frequency of the corresponding mode of the RS, $\omega_{I}$, $\omega_{P}$, and $\omega_{S}$ for the idler, pump, and signal CROWS respectively. The resultant single-defect cavities chosen have radii of 66, 76, and 89 nm respectively for the idler, pump, and signal CROWs. All three CROWs have a period $d=3a$. The dispersion curves for the three CROWs are shown in Fig. 2.
Figure 2: Dispersion for the CROW bands as calculated in the nearest-neighbor tight-binding approximation (see Sec. III.1.3). Idler, pump, and signal CROW bands are shown in red, green, and blue, respectively, with corresponding RS QM frequencies displayed as dashed lines intersecting their corresponding CROW band.
Note that the resonant frequencies are at the center of each CROW band and that $\omega_{S}$ and $\omega_{I}$ lie outside the pump band. Thus, the pump light does not contaminate the output light in the two other CROWS. As we shall discuss in Section III, the generated light will exit the resonant structure via the signal and idler CROWs. Because the signal and idler bands do not overlap, there will be no cross-contamination of the signal and idler modes at the output. The band-pass filtering functionality is one of the reasons that we have used CROWS rather than line-defect waveguides to couple the light into and out of the RS. The key properties of the system are summarized in Table 1.
@@
CROW Cav. Rad. [nm] 76 89 66
CROW Ctr. Freq. [THz] 192.3 193.2 191.2
Coupling Cav. Rad. [nm] - 91 55
@@ III Theoretical Approach
In this section, we present our theoretical approach to simulating the generation of quantum states of light via nonlinear processes in a PC CCS.
Our approach can be broken into four distinct steps, shown pictorially in Fig. 3. Before presenting detailed discussions of each of the steps, we start with an overview of each of the steps.

### v3

@@ Abstract
Entangled multi-mode squeezed states of light have a wide variety of applications in quantum information systems, particularly in the generation of non-Gaussian states of light, which are central to continuous-variable quantum computing. Although theoretical approaches exist to model the nonlinear generation of one- and two-mode entangled states of light in ring resonator systems, these approaches are difficult to implement in modeling more complicated many-cavity systems. In this work, we present an efficient and accurate theoretical approach to modeling the generation and propagation of quantum states of light in lossy coupled-cavity systems containing a two- or three-mode nonlinear resonant structure. Our approach is general and computationally viable even in systems with hundreds of modes. We apply our method to the design and modeling of a multimode photonic crystal coupled-cavity system for the generation of entangled squeezed states of light on-chip. The system consists of a three-mode resonant structure coupled to three coupled-resonator optical waveguides (CROWs) in a square lattice silicon photonic crystal slab. The computational speed of the method allows us to efficiently optimize the system such that the signal and idler light in the two output CROWs remains entangled even after propagating tens of cavities down the CROWs.
@@ I Introduction
Entangled squeezed thermal states of light [Pfister2020, Adesso2014] are sought for their entanglement and noise properties. They have applications in continuous-variable quantum computing [Alexander2024, Feldman2024, Menicucci2006], quantum key distribution [Leverrier2011], quantum-enhanced sensing [Dorfman2016, LIGO, Schnabel2017], and the generation of non-Gaussian entangled states of light [Tomoda2024]. Consequently, there is considerable interest in modeling, designing, and optimizing nanophotonic platforms for the generation of entangled quantum states of light.
@@
Entangled, multi-mode, squeezed thermal states of light may be created by generating single-mode squeezed thermal states of light in an optical cavity [Seifoory2017, Vendromin2020] and interfering them using a beam-splitter array to build multi-mode entanglement [Madsen2022]. Alternatively, they can be generated directly via a nonlinear process described by a Hamiltonian containing terms with an operator of the form $(a^{\dagger}_{p}a^{\dagger}_{q})^{n},\ n\in\mathbb{N}^{>0}$, where $a_{p}^{\dagger}$ is the creation operator for photons in the $p^{th}$ mode.
Nonlinear devices using microring resonators have been explored both experimentally [Silverstone2015, Tan2019, Zhang2021, Madsen2022] and theoretically [Quesada2021, Mataji-Kojouri2024, Vendromin2020, Vendromin2021, Vendromin2022, Vendromin2023, Vendromin2024] for the generation of entangled squeezed states of light. Microring resonator systems can be fabricated with very high quality factors (Q) [Zhang2021] and are relatively simple to design, fabricate, and couple to channel waveguides. Moreover, due to the relative simplicity of the ring modes as well as their coupling to the channel waveguide, accurate modeling of state generation in these systems can be achieved using only a few ring modes along with a well-established input-output formalism [Zhang2021, Dignam2025]. There are, however, two important drawbacks of ring resonator systems. First, they are relatively large, with typical ring radii of 50 to 100 microns. Second, the rings support a large number of nearly equally spaced modes, which can lead to parasitic nonlinear processes that degrade the properties of the desired quantum state [Zhang2021, Seifoory2022] unless more complicated coupled-ring systems are employed [Zhang2021] [Larsen2025]. It is therefore salient to explore alternative platforms for multimode squeezed state generation, aiming to improve the footprint and optical properties.
One alternative to ring resonator systems are photonic crystal (PC) coupled cavity systems (CCSs). Such systems have been employed for the generation of photon pairs [Azzini2013] and have been investigated theoretically [KamandarDezfouli2017, Seifoory2018, Seifoory2019] for the generation of squeezed states of light. Although the quality factors of the resonant modes in PC CCSs are generally an order of magnitude or more smaller than that of ring resonator systems, the resonant cavities (and mode volumes) are much smaller, with cavity sizes on the order of hundreds of nanometers. Moreover, they can be designed such that they only support the modes required for the desired nonlinear optical processes. Finally, employing coupled resonator optical waveguides (CROWs) for the output ports, signal filtering is integrated in the design, removing the need to include additional filters, which can add considerable loss [Larsen2025].
@@
In this work, we examine PC CCSs as an alternative platform to microring resonators for the generation of squeezed states (see Fig. 1). The system consists of a three-mode, three-defect resonant structure (RS) coupled to three CROWs [Yariv1999] in a square lattice silicon photonic crystal slab (see Fig. 1). The key element of the system is the RS: it is comprised of three single-defect cavities and supports three resonant modes, one each at the signal, pump, and idler frequencies. The RS enhances the nonlinear spontaneous four wave mixing (SFWM) process that generates the desired entangled state. An entangled two-mode squeezed thermal state (2MSTS) is generated in the RS at the signal and idler frequencies via SFWM when excited by a Gaussian pump pulse with a frequency that is mid-way between the signal and idler frequencies. The generated light propagates out of the RS as two entangled pulses at the signal and idler frequencies in spatially-distinct CROWs that are designed to block any pump light.
Any model of squeezed state generation in nanophotonic systems must include: 1) the nonlinear generation process; 2) the effects of loss; and 3) the coupling of the generated light into the output channel waveguide(s). In ring resonators, because the mode structure of the combined ring and channel waveguide system is relatively simple, with the loss being essentially the same for all modes, the nonlinear evolution and mode propagation can be treated accurately using an input-output formalism, with the loss accounted for either using phantom channels [Quesada2021], or using a density matrix formalism for the ring modes [Dignam2025]. However, in PC coupled cavity systems, the mode structure can be complicated, with the loss being very different in different modes, and with convoluted pump and signal propagation dynamics. Although one could, in principle, model such systems using a multi-mode density matrix formalism, such an approach quickly becomes computationally intractable because the number of Fock basis states required for convergence grows as $(N+1)^{M}$, where $M$ is the number of modes and $N$ is the maximum number of photons that may be generated in each mode. For example, in this work we study the generation and propagation of a squeezed thermal state in a system of 184 coupled cavities; modeling this state of light in such a system would require a minimum of $M=33$ modes  [^1] We arrive at the minimum mode number of $M=33$ by projecting the state generated in the RS onto the full-system QMs, and counting the number of full-system QMs with at least $1\%$ of the total number of photons.. If we permit each of these 33 modes to have up to $N=2$ photons, then we require $3^{33}$, or more than $5\times 10^{15}$ total Fock basis states. Such a problem is clearly intractable when working in a Fock basis.
In this work, we develop a new theoretical approach to simulate the nonlinear generation of entangled multi-mode squeezed thermal states of light in systems of coupled lossy nanophotonic resonators in which there is a localized region – the RS – where the nonlinear interaction occurs. Our approach enables the simulation and study of quantum states of light in a complex PC CCS that cannot easily be modeled using traditional Fock-state based techniques.
We apply our approach to the simulation and optimization of the PC CCS shown in Fig. 1, but note that our method can be applied to many other nanophotonic resonator systems. The simulation is carried out in four distinct steps, as shown in Fig. 2. First, we determine the lossy quasi-modes (QMs) of the full system under consideration. To accomplish this, we use finite difference time domain (FDTD) simulations to compute the electric field profile and quality factor of the QMs of the single-defect cavities from which we construct the PC CCS [Fussell2007]. We then use a tight-binding approach to construct the QMs of the full system from these single-defect modes [Fussell2007, Fussell2008, Dignam2012, KamandarDezfouli2014]. Second, we use this basis of system QMs to compute the classical evolution of the pump light (for a given initial pump field) and determine the pumping-field amplitude inside the RS as a function of time. Third, we compute the time evolution of density operator of the quantum state of light generated via nonlinear interaction in the RS. Fourth, we determine the quantum evolution of the generated state throughout the system by solving the Adjoint master equation after the pump has largely left the RS.
Our simulations show that strong squeezing can be obtained in the RS and that entanglement between the light and the signal CROWs persists for many tens of cavities down the waveguides. However, we find that entanglement between the signal and idler is significantly degraded in the CROWs. We analyze the source of this degradation and point to possible solutions.
@@
The paper is organized as follows: in Section II, we provide a detailed description of the designed device and outline the design goals; in Section III, we introduce the four-step analysis procedure we developed to simulate the designed system; in Section IV, we present and discuss the results of the analysis; and finally in Section V, we conclude.
@@ II System Overview and Design Goals
In this section, we describe the PC CCS that we have designed and optimized using our theoretical approach. We present our design process, outline the objectives of our design, and present the specifications of the designed PC CCS within which we will simulate the generation and propagation of multimode squeezed light. The details of the calculation of the coupled cavity modes are presented in Appendix A.
The PC CCS consists of a three-mode RS and three CROWs in a PC slab, shown schematically in Fig. 1. The underlying PC slab is a 384 nm thick silicon (n=3.47) slab with air holes of radius 192 nm etched in a rectangular lattice with a lattice constant $a=480$ nm. This particular square lattice geometry of the PC was chosen because the single-defect cavities formed in this PC support a single relatively high-Q mode in the photonic band-gap of the PC [Fussell2007].
We have designed the PC CCS to generate and propagate an entangled 2M STS with high brightness and robust entanglement. We have four main objectives for the design of the PC CCS. First, we need a RS with spatially-overlapping modes at the frequencies of the pump ($\omega_{P}$), signal ($\omega_{S}$), and idler ($\omega_{I}$), where $2\omega_{p}=\omega_{s}+\omega_{I}$ to enable SFWM between the pump, signal, and idler modes. Second, we need to be able to inject the pump light into the RS via a CROW such that there is a large pump intensity buildup inside of the RS. Third, we require that the quantum state of light generated in the RS can couple out of the RS into two separate CROWs, with central frequencies $\omega_{s}$ and $\omega_{I}$ that do not allow propagation of the pump light. The key properties of the system are summarized in Table 1.
(a) Idler, $\tilde{\omega}_{I}=0.30610-i1.19\times 10^{-5}\ (2\pi c/a)$
(b) Pump, $\tilde{\omega}_{P}=0.30763-i8.17\times 10^{-6}\ (2\pi c/a)$
(c) Signal, $\tilde{\omega}_{S}=0.30916-i3.69\times 10^{-6}\ (2\pi c/a)$
Figure 3: Plot of $|C_{j,x}(x,y,0)|^{2}$ for the three RS QMs. The area of each subfigure is enclosed by the purple rectangle in Fig. 1. The complex frequency for each of the QMs (when the CROWs are absent) are given below each plot.
@@
Figure 4: Dispersion for the CROW bands as calculated in the nearest-neighbor tight-binding approximation (see Appendix A). Idler, pump, and signal CROW bands are shown in red, green, and blue, respectively, with corresponding RS QM frequencies displayed as dashed lines intersecting their corresponding CROW band.
We next designed the three CROW structures used to couple the pump light into and generated light out of the RS. The dispersion curves for the three CROWs are shown in Fig. 4. The CROWS were designed so that the center frequency of each band matches the corresponding RS QM frequency (see Table 1) and their bandwidths encompass only one of the pump, signal, or idler frequencies. This was achieved by constructing each CROW from a set of single-defect cavities with a resonant frequency equal to the frequency of the corresponding mode of the RS, and by setting the period to $d=3a$ (for all three CROWs). The resultant single-defect cavities chosen have radii of 66, 76, and 89 nm respectively for the idler, pump, and signal CROWs. Because the signal and idler CROWs can’t pass light at the pump frequency, there will be no pump contamination in the output CROWS. Moreover, because the pass bands of signal and idler CROWs do not overlap, the light exiting in the signal and idler CROWs will have no frequency overlap.
@@
CROW Cav. Rad. (nm) 76 89 66
CROW Ctr. Freq. (THz) 192.3 193.2 191.2
Coupling Cav. Rad. (nm) - 91 55
@@ III Theoretical Approach
In this section, we present our theoretical approach to simulating the generation of quantum states of light via nonlinear processes in a PC CCS. The subsections describe the following four steps that we employ to simulate the generation and time evolution of the quantum state of light (see Fig. 2): A) Calculation of the quasimode basis states; B) Calculation of the classical pump evolution; C) Modeling the nonlinear generation of the quantum state in the RS; and D) Calculation of the free evolution of the quantum state of light in the full PC CCS. This section provides an overview of our approach; the details can be found in the appendices cited in the text.
III.1 Quasi-Mode Bases